OpenStudy (anonymous):

F(x)=2x+5; g(x)=5x-4 find f(g)

OpenStudy (anonymous):

u mean f(g(x))?

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

\[f(g(x))=f(5x-4)=2(5x-4)+5\] you can simplify the last one a bit

OpenStudy (anonymous):

Ok then what?

OpenStudy (anonymous):

If that is f(g(x)) ... This is how to solve... F(x)=2x+5; g(x)=5x-4 find f(g(x)) f(g(x))=2(5x-4)+5 = 10x-8+5=10x-3

OpenStudy (anonymous):

Thus, f(g(x)) of F(x)=2x+5; g(x)=5x-4 is 10x-3

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

no problem...